OpenStudy (ellecullen):

Homework help ASAP please: 5. According to Keynesian theory, what is the most important determinant of households’ spending on goods and services? A. The price level B. The interest rate C. Autonomous consumption D. Disposable income E. Government spending 9. At the point where the disposable income line intersects the consumption function, saving: A. equals consumption. B. equals disposable income. C. is less than zero. D. is equal to zero. E. is greater than consumption. The rest is in comments

OpenStudy (ellecullen):

@bohotness

OpenStudy (bohotness):

which one do you think is you answr love

OpenStudy (ellecullen):

5, I don't know, maybe D. 9. E maybe. What do you think?

OpenStudy (bohotness):

i agree love

OpenStudy (ellecullen):

1. Consider an economy in which C = 1000 + 0.75YD, I = 200, G = 300, T = 300, and (X-M) = 0. A. Calculate equilibrium GDP. (3 points) B. Calculate saving at this level of GDP. Hint: Use the equation for leakages. (4 points) C. Calculate the spending multiplier. (2 points) D. Calculate the tax multiplier. (2 points) E. How will equilibrium GDP change if G increases by 100, and what is the new equilibrium level of GDP? F. How will equilibrium GDP change if G and T both increase by 500, and what is the new equilibrium level of GDP?

OpenStudy (bohotness):

what do you tihkn your answer is love

OpenStudy (ellecullen):

I need help on these

OpenStudy (ellecullen):

I don't know. Can you help with this?

OpenStudy (bohotness):

i will try

OpenStudy (ellecullen):

Formulas I have Y = C + I + G+ (X – M) * = times Y = [1 / ( 1 – b) * ( a + I + G + NX) ] – [ b / (1 – b) * T]

OpenStudy (bohotness):

thats right

OpenStudy (ellecullen):

A. Y = [ 1 /(1 – 0.75) * (1,000+ 200 + 300 +0)] – [ 0.75 / (1-0.75)] Y = [ 4 * 1500] – [ 3 * 300] Y = [6000] – [900] = 5100

OpenStudy (bohotness):

XD

OpenStudy (ellecullen):

leakages equation S+ T+M

OpenStudy (ellecullen):

Can you help calculate this please?

OpenStudy (ellecullen):

@bohotness, i'm having trouble calculating, as i don't know what S is .

OpenStudy (bohotness):

hold on i am ba at that

OpenStudy (ellecullen):

A. Calculate the spending multiplier. (2 points) 1 / (1 –b)  1 / ( 1 - 0.75) = 4 Spending Multiplier = 4

OpenStudy (ellecullen):

sorry, spending multiplier is C. part

OpenStudy (bohotness):

okay and its okay

OpenStudy (ellecullen):

I am having trouble with B. Any ideas?

OpenStudy (ellecullen):

D. Calculate the tax multiplier. (2 points) MPC / ( 1 – MPC) MPC = C / Yd MPC = 0.208, rounded to 0.21 Yd=Y – T  4800 0.21 / ( 1 -0.21) = 0.3 rounded from 0.26

OpenStudy (bohotness):

okay i thik you got d

OpenStudy (ellecullen):

E. How will equilibrium GDP change if G increases by 100, and what is the new equilibrium level of GDP? (4 points) Y = [ 1 / ( 1- 0.75) * ( 1000 + 200 + 400+ 0)] – [0.75 / ( 1 - 0.75) * T ] Y = [ 4 * 1600 ] – [ 3 * 300] Y = [ 6400 ] – [ 900 ] Y = 5500 The equilibrium GDP will rise from 5100 to 5500,the new equilibrium level of GDP is that Y = 5500

OpenStudy (bohotness):

XDD okay you doing it can you show be b?

OpenStudy (ellecullen):

F. How will equilibrium GDP change if G and T both increase by 500, and what is the new equilibrium level of GDP? (3 points) Y = [ 1 / ( 1- 0.75) * ( 1000 + 200 + 800 + 0) ] – [0.75 / ( 1 – 0.75) * 800] Y = [ 4 * 2000 ] – [ 3 * 800 ] Y = [ 8000 ] – [ 2400 ] Y = 5600, or rounded to 6000 The equilibrium GDP will increase again, and with Y being 5600 or rounded, the new equilibrium level is at 6000.

OpenStudy (ellecullen):

b. I only have that leakages equation id S + T + M For M, I only know that (x - M) = 0. T = 300. But I do not know the S. This is all I have on b.

OpenStudy (bohotness):

wait you are on the right true love